OpenStudy (anonymous):

anita has 2 sisters and 3 brothers. the mean of all their ages is 6 yrs. what will their mean age be 10 yrs from now? 20 yrs from now?

OpenStudy (anonymous):

assume a,b,c,d,e are the ages of 2 sisters and 3 brothers a+b+c+d+e=6*5 a+b+c+d+e+x=10*5 =>x=20=>4 years a+b+c+d+e+y=20*5=>y=70=>14 years

OpenStudy (anonymous):

first work out the age that they'll be then try to work out the mean....... actually you can't do that because ?! well there is some thing to do with dividing! I think!

OpenStudy (dumbcow):

fyi minnie i answered your question on the chat

OpenStudy (anonymous):

yeah- thanks for that!